The demand for comic books is given by P = 90 – Qd. Supply of comic books is given by P = 2Qs. Suppose that the government imposed a price floor of $70 in this market. What will be the minimum size of deadweight loss?
1
Expert's answer
2015-04-04T11:53:13-0400
P = 90 – Qd, P = 2Qs
In this case equilibrium quantity is:
90 - Q = 2Q
Q = 30 units
Equilibrium price is:
P = 2*30 = $60
If the government imposes a price floor of $70 in this market, there will be a deadweight loss.
New quantity demanded is Qd = 90 - 70 = 20 units.
New quantity supplied is Qs = 70/2 = 35 units.
The minimum size of deadweight loss is DWL = 0.5*(35 - 20)*(70 - 60) = $75